OpenStudy (hang254):

Geometry Help! The vertices of triangle ABC on a coordinate grid are A(-1, -1), B(-5, -5), and C(-1, -5). The vertices of triangle AQR are A(-1, -1), Q(-3, -3), and R(-1, -3). Which proportion can be used as a step in proving that triangle ABC is similar to triangle AQR?

OpenStudy (hang254):

Possible answers: ab/qb=2√8/√8 bc/aq=4/4√8 bc/qr=4/4 ab/aq=2√8/√8

OpenStudy (hang254):

@jim_thompson5910 can you help me with this problem?

jimthompson5910 (jim_thompson5910):

The side AB corresponds to AQ (since these are the first two points). The side BC corresponds to QR (these are the last two). And finally, CA or AC corresponds to RA or AR (first and last points) Now you have to find the distances between each point. Once you know the distances, you can use them with the corresponding sides to set up a ratio. This will prove that the two triangles are similar.

OpenStudy (hang254):

ok, thank you

jimthompson5910 (jim_thompson5910):

yw